OpenStudy (anonymous):

Solve: –4(2x – 1) = –6(x + 2) – 2

OpenStudy (anonymous):

the first thing, would be to multiply, the number outside each parenthesis, to the inside, using the distributive property. -4(2x - 1) = -6(x + 2) - 2 -4(2x) -4(-1) = -6(x) -6(2) - 2 -8x + 4 = -6x - 12 - 2 -8x + 4 = -6x - 14 then we rearrange it so the common ones can be together 4 = -6x + 8x - 14 4 = 2x - 14 4 + 14 = 2x 18 = 2x then we pass the two, by dividing since it was multiplying 18/2 = x 9 =x x = 9, that would be your answer